OpenStudy (anonymous):

please help with math I will fan u find the coordinates of the midpoint whose endpoint are A (-3,-3) and B (6,6)

OpenStudy (anonymous):

step by step n explain what u did please

OpenStudy (imstuck):

Ok, I would love to do that for you. There is a midpoint formula we will use to find what you are looking for. It is\[(\frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 })\]What you do is add the x coordinates from both the points and divide the sum by 2, then you add the y coordinates from both the points and divide that sum by 2, and you get one new point that is the midpoint. Here you have x coordinates of 6 and -3. When you add those you get 3. Divide 3 by 2 and the x coordinate of your midpoint is 3/2. Now do the same with the y coordinates. They are also -3 and 6. So when you add those and divide by 2, you get 3/2 there too. So the coordinate of your midpoint is \[(\frac{ 3 }{ 2 },\frac{ 3 }{ 2 })\]Any other ones you would like help on?

jagr2713 (jagr2713):

yea what stuck said, use midpoint formula and plug the numbers into the formula